The equations x + 1 = 0 and 2x + 2 = 0 have the same solution. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.
Consider equations x + 1 = 0

So,
x = -1
Consider the equation, 2x + 2 = 0
Divide both the side by 2,
Then we get
x + 1 = 0
Therefore,
x = - 1

Thanks